Theory and Applications of Categories, Vol. 30, No. 53, 2015, pp. 1841–1863.

PROFINITE TOPOLOGICAL SPACES

G. BEZHANISHVILI, D. GABELAIA, M. JIBLADZE, P. J. MORANDI

Abstract. It is well known [Hoc69, Joy71] that profinite T_{0}-spaces are exactly the spectral spaces. We generalize this result to the category of all topological spaces by showing that the following conditions are equivalent:

1 X, τ is a profinite topological space.

2 The T_{0}-reflection of X, τ is a profinite T_{0}-space.

3 X, τ is a quasi spectral space (in the sense of [BMM08]).

4 X, τ admits a stronger Stone topology π such that X, τ, π is a bitopological quasi spectral space (see Definition 6.1).

1. Introduction

A topological space is profinite if it is (homeomorphic to) the inverse limit of an inverse system of finite topological spaces. It is well known [Hoc69, Joy71] that profinite T_{0}- spaces are exactly the spectral spaces. This can be seen as follows. A direct calculation shows that the inverse limit of an inverse system of finite T_{0}-spaces is spectral. Conversely, by [Cor75], the category Spec of spectral spaces and spectral maps is isomorphic to the category Pries of Priestley spaces and continuous order preserving maps. This isomorphism is a restriction of a more general isomorphism between the category StKSp of stably compact spaces and proper maps and the category Nach of Nachbin spaces and continuous order preserving maps [GHKLMS03]. Priestley spaces are exactly the profinite objects in Nach , and the proof of this fact is a straightforward generalization of the proof that Stone spaces are profinite objects in the category of compact Hausdorff spaces and continuous maps. Consequently, spectral spaces are profinite objects in StKSp . From this it follows that spectral spaces are profinite objects in the category of T_{0}-spaces and continuous maps.

We aim to generalize these results to characterize all profinite topological spaces. For this we first note that Nachbin spaces are partially ordered compact spaces such that the order is closed. It is natural, as Nachbin himself did in [Nac65], to work more generally with preordered Nachbin spaces. As it follows from [Dia07, Sec. 2] (see also [JS08, Sec. 1], and more generally, [Hof02, Sec. 4]), the fact that Priestley spaces are exactly the profinite objects in Nach generalizes to the setting of preordered spaces. Consequently, preordered Priestley spaces are exactly the profinite objects in the category PNach of preordered Nachbin spaces. Associating with each preordered Nachbin space the topology of open upsets defines a functor K from PNach to the category Top of all topological spaces and continuous maps. It follows from [HST14, Sec. III.5.3] that K PNach Top has a left adjoint M Top PNach . From this we derive that a topological space is profinite iff it is the K-image of a preordered Priestley space.

We are very thankful to the referee for many useful suggestions and pointers to the literature.

Received by the editors 2014-11-04 and, in revised form, 2015-11-19.

Transmitted by Walter Tholen. Published on 2015-12-03.

2010 Mathematics Subject Classification: 18B30, 18A30, 54E55, 54F05, 06E15.

Key words and phrases: Profinite space, spectral space, stably compact space, bitopological space, ordered topological space, Priestley space.

© G. Bezhanishvili, D. Gabelaia, M. Jibladze, P. J. Morandi, 2015. Permission to copy for private use granted.

To characterize the K-image of PNach in Top, note that the isomorphism between StKSp and Nach generalizes as follows. Dropping the T_{0} separation axiom from the definition of a stably compact space results in the concept of a quasi stably compact space, which generalizes the concept of a quasi spectral space of [BMM08], and corresponds to that of a representable space of [HST14, Sec. III.5.7]. However, unlike the case with stably compact spaces, the patch topology of a quasi stably compact space may not be compact Hausdorff. In order to obtain a category isomorphic to PNach , we have to introduce bitopological analogues of quasi stably compact spaces, which together with the quasi stably compact topology also carry a stronger compact Hausdorff topology. The same way PNach is isomorphic to the category of Eilenberg-Moore algebras for the ultrafilter monad on the category of preordered sets [Tho09], the category BQStKSp of bitopological quasi stably compact spaces is isomorphic to the category of Eilenberg-Moore algebras for the ultrafilter monad on Top. Consequently, applying [HST14, Cor. III.5.4.2] yields that PNach is isomorphic to BQStKSp , and this isomorphism restricts to an isomorphism of the category PPries of preordered Priestley spaces and the category BQSpec of bitopological quasi spectral spaces. This characterizes the spaces in the K-image of PNach as the bitopological quasi stably compact spaces, and the spaces in the K-image of PPries as the bitopological quasi spectral spaces. Thus, profinite topological spaces are exactly the bitopological quasi spectral spaces.

Our main result establishes that each quasi stably compact space X, τ can be made into a bitopological quasi stably compact space X, τ, π , which will become a bitopological quasi spectral space if X, τ is quasi spectral. Since a space is quasi spectral iff its T_{0}-reflection is spectral, we obtain that a topological space is profinite iff so is its T_{0}-reflection. This yields that for a topological space X, τ , the following conditions are equivalent:

1 X, τ is a profinite topological space.

2 The T_{0}-reflection of X, τ is a profinite T_{0}-space.

3 X, τ is a quasi spectral space.

4 X, τ admits a stronger Stone topology π such that X, τ, π is a bitopological quasi spectral space.

2. Preliminaries

In this section we recall the basic topological notions that are used in the paper. We start by the following well-known definition.

2.1. Definition. Let X be a topological space.

1 A subset A of X is saturated if it is an intersection of open subsets of X, and it is irreducible if from A F G, with F and G closed, it follows that A F or A G.

2 The space X is sober if each irreducible closed set of X is the closure of a point of X.

3 The space X is stable if compact saturated subsets of X are closed under finite intersections, and coherent if compact open subsets of X are closed under finite intersections and form a basis for the topology.

4 The space X is locally compact if for each x X and each open neighborhood U of x, there exist an open set V and a compact set K such that x V K U.

5 The space X is stably compact if X is a compact, locally compact, stable, sober T_{0}-space.

6 The space X is spectral if X is a compact, coherent, sober T_{0}-space.

2.2. Remark. It is customary to call a space X sober if each irreducible closed set is the closure of a unique point. This stronger definition of soberness automatically implies that the space X is T_{0}. Since we will work with non-T_{0}-spaces, we prefer the weaker definition of soberness given above, and will explicitly add the T_{0} separation axiom when needed. In [HST14, Sec. III.5.6] this weaker condition is referred to as weakly sober.

It is well known that each spectral space is stably compact. In fact, stably compact spaces are exactly the retracts of spectral spaces; see [Joh82, Thm. VII.4.6] or [Sim82, Lem. 3.13].

2.3. Definition.

1 A continuous map f X X^{}between stably compact spaces is proper if the f-inverse image of a compact saturated subset of X^{}is compact in X.

2 A continuous map f X X^{}between spectral spaces is spectral if the f-inverse image of a compact open subset of X^{}is compact in X.

Let StKSp be the category of stably compact spaces and proper maps, and let Spec be the category of spectral spaces and spectral maps. It is well known that a continuous map f X X^{}between spectral spaces is spectral iff it is proper. Thus, Spec is a full subcategory of StKSp . In fact, Spec can be characterized as the profinite objects in StKSp or as the profinite objects in the category of T_{0}-spaces (see [Hoc69, Prop. 10], [Joy71]).

2.4. Definition.

1 A preordered space is a triple X, π, , where π is a topology and is a preorder on X (reflexive and transitive relation on X). If is a partial order, then X, π, is an ordered space.

2 The preordered space X, π, is a preordered Nachbin space if X, π is compact Hausdorff and is closed in the product X X. If is a partial order, then X, π, is a Nachbin space.

3 A subset U of X is an upset provided x U and x y imply y U. The concept of a downset is defined dually.

4 The preordered space X, π, satisfies the Priestley separation axiom if from x y it follows that there is a clopen upset U containing x and missing y.

5 The preordered space X, π, is a preordered Priestley space if X, π is a Stone space (zero-dimensional compact Hausdorff space) and X, π, satisfies the Priestley separation axiom. If is a partial order, then X, π, is a Priestley space.

Let X, π be a compact Hausdorff space. It is well known (see, e.g., [Nac65, Sec. I.1 and I.3] or [BMM02, Prop. 2.3]) that a preorder on X is closed iff for all x, y X, from x y it follows that there exist an open upset U and an open downset V such that x U, y V, and U V . From this it is clear that each preordered Priestley space is a preordered Nachbin space.

Let PNach be the category of preordered Nachbin spaces and continuous preorder preserving maps (x y implies f x f y ), and let Nach be its full subcategory consisting of Nachbin spaces. Let also PPries be the full subcategory of PNach consisting of preordered Priestley spaces, and Pries be the full subcategory of PPries consisting of Priestley spaces. It is well known (see, e.g., [Spe72]) that the objects of Pries are exactly the profinite objects of Nach . This result generalizes to the setting of preordered spaces. Indeed, as follows from [Dia07, Sec. 2] (see also [JS08, Sec. 1], and more generally, [Hof02, Sec. 4]), the objects of PPries are exactly the profinite objects of PNach .

2.5. Remark. That Priestley spaces are exactly the profinite objects of Nach can be proved directly or by using Priestley duality [Pri70], according to which Pries is dually equivalent to the category Dist of bounded distributive lattices and bounded lattice homomorphisms. This dual equivalence restricts to the dual equivalence of finite distributive lattices and finite posets. Since Dist is locally finite (each finitely generated object in Dist is finite), each bounded distributive lattice is the direct limit of a direct system of finite distributive lattices. This, by Priestley duality, yields that each Priestley space is the inverse limit of an inverse system of finite posets. Consequently, Priestley spaces are exactly the profinite objects of Nach .

That preordered Priestley spaces are exactly the profinite objects of PNach also has a similar alternate proof. As follows from [Bez13, Thm. 5.2], PPries is dually equivalent to the category BDA of pairs B, D , where B is a Boolean algebra and D is a bounded sublattice of B. This dual equivalence restricts to the dual equivalence of finite preorders and finite objects of BDA . Since each Boolean algebra and each distributive lattice is the direct limit of a direct system of their finite subobjects, it follows that each B, D BDA is also the direct limit of its finite subobjects in BDA . Therefore, by the duality of [Bez13], each preordered Priestley space is the inverse limit of an inverse system of finite preorders. Thus, preordered Priestley spaces are exactly the profinite objects of PNach .

The category of stably compact spaces is isomorphic to the category of Nachbin spaces [GHKLMS03], and this isomorphism restricts to an isomorphism between the categories of spectral spaces and Priestley spaces [Cor75]. It is obtained as follows. Let X, τ be a stably compact space. The co-compact topology τ^{k} is the topology having compact saturated sets as closed sets, and the patch topology π is the smallest topology containing τ and τ^{k}; that is, π τ τ^{k}. The specialization order of τ is given by x y iff x cl_{τ} y . This is a partial order because X, τ is a T_{0}-space. Then X, π, turns out to be a Nachbin space. Moreover, a map f X X^{}between stably compact spaces X, τ and X^{}, τ^{}is proper iff it is continuous and order preserving between the corresponding Nachbin spaces X, π, and X^{}, π^{}, ^{}. This defines a functor F StKSp Nach .

Conversely, if X, π, is a Nachbin space, then let π^{u} be the topology of open upsets (and π^{d} be the topology of open downsets). Then X, π^{u} is a stably compact space (and so is X, π^{d} ). Moreover, a map f X X^{}between Nachbin spaces X, π, and X^{}, π^{}, ^{}is continuous order preserving iff it is a proper map between the corresponding stably compact spaces X, π^{u} and X^{}, π^{}^{u} . This defines a functor G Nach StKSp . The functors F, G establish the desired isomorphism between StKSp and Nach .

This isomorphism restricts to an isomorphism between Spec and Pries . In fact, if X, τ is a spectral space, then the patch topology can alternatively be defined as the topology generated by compact opens of X, τ and their complements (which are compact opens in the spectral space X, τ^{k} ). As a result, we arrive at the following commutative diagram, where the horizontal arrows indicate an isomorphism of categories, while the vertical ones indicate that one category is a full subcategory of another.

StKSp  Nach

Spec  Pries

We will see later on how to generalize this commutative diagram to involve PNach and PPries .

3. Profinite topological spaces

Let X, π, PNach and let τ π^{u} be the topology of open upsets. Then X, τ Top. Moreover, if f X, π, X^{}, π^{}, ^{}is continuous and preorder preserving, then f X, τ X^{}, τ^{}is continuous. Thus, we have a functor K PNach Top. It follows from general considerations in [HST14, Cor. III.5.3.4] that K PNach Top has a left adjoint M Top PNach . The description of M can be derived from [HST14, Thm. III.5.3.5 and Ex. III.5.3.7(1)] and is closely related to Salbany’s construction [Sal00] (see also [BMM08]). More precisely, let β Set Set be the ultrafilter functor that assigns to a set X the set βX of ultrafilters on X, and to a map f X Y the map βf βX βY given by

βf χ T Y f^{1} T χ .

For X, τ Top, let M X, τ βX, Π, . Here Π is the Stone topology on βX given by the basis φ S S X , where φ X βX is the Stone map

φ S χ βX S χ .

Moreover, for χ, ξ βX, we set χ ξ iff χ τ ξ. It is a consequence of [BMM08, Thm. 3.8] that βX, Π, PPries PNach .

For a map f X X^{}, we have βf ^{1}φ S φ f^{1}S , and so βf βX, Π βX^{}, Π^{}is continuous. Furthermore, if f X, τ X^{}, τ^{}is continuous, then χ ξ implies βf χ ^{}βf ξ . Therefore, M Top PNach is a functor.

3.1. Proposition. [HST14, Sec. III.5] The functor M is left adjoint to K.

We are ready to prove that profinite topological spaces are exactly the K-images of PPries .

3.2. Theorem. A topological space X, τ is profinite iff X, τ is homeomorphic to an object of K PPries .

Proof. First suppose X, π, PPries . By [Dia07, Sec. 2], X, π, is an inverse limit of a diagram of finite objects of PNach . Since K has a left adjoint, K preserves limits. Therefore, K preserves the limit of this diagram. As the values of K on finite objects in PNach are finite topological spaces, we conclude that K X, π, is a profinite topological space.

Conversely, suppose X, τ is a profinite topological space. Then X, τ is the inverse limit of a diagram of finite topological spaces in Top. Applying M to produces a diagram ˜ of finite objects in PNach . Obviously K ˜ . Let X̃, π̃, ˜ be the inverse limit of ˜ in PNach . By [Dia07, Sec. 2], X̃, π̃, ˜ PPries . Since K preserves the limit of ˜ , we have K X̃, π̃, ˜ K \begin{matrix}lim \\ \end{matrix}˜\begin{matrix}lim \\ \end{matrix}K ˜ \begin{matrix}lim \\ \end{matrix}X, τ . Thus, X, τ is homeomorphic to an object of K PPries .

In Section 6 we characterize K PPries within Top. This requires preparation, which is the subject of the next section.

4. Algebras for the ultrafilter monad

We recall that a monad T T, e, m on a category C consists of an endofunctor T C C together with natural transformations e Id_{C} T (unit) and m T T T (multiplication) satisfying m T e m eT Id_{T} and m T m m mT.

\begin{matrix}T & T T \\ T T & T \end{matrix} ; T e ; eT ; m ; m  \begin{matrix}T T T & T T \\ T T & T \end{matrix} ; T m ; mT ; m ; m

A T-algebra or an Eilenberg-Moore algebra is a pair X, a , where X is a C-object and a T X X is a C-morphism satisfying a e_{X} Id_{X} and a T a a m_{X}.

\begin{aligned}X & T X \\ & X \end{aligned} ; e_{X} ; a  \begin{matrix}T T X & T X \\ T X & X \end{matrix} ; T a ; m_{X} ; a ; a

A T-homomorphism f X^{}, a^{}X, a is a C-morphism f X^{}X satisfying a T f f a^{}.

\begin{matrix}T X^{} & T X \\ X^{} & X \end{matrix} T f a^{}a f

It is easy to see that T-algebras form a category, which we denote by C^{T}.

We are mainly interested in the ultrafilter monad βββ β, e, m . The unit e is given by the embeddings e_{X} X βX assigning to x X the principal ultrafilter e_{X} x χ_{x}; and the multiplication m consists of the maps m_{X} ββX βX given by

m_{X} Ξ S X φ S Ξ ,

where Ξ ββX and φ is the Stone map.

By Manes’ theorem [Man69] (see also [HST14, Thm. III.2.3.3]), the category of βββalgebras is isomorphic to the category of compact Hausdorff spaces and continuous maps. Under this isomorphism, a compact Hausdorff space X, π corresponds to the βββ-algebra X, lim_{π} , where the map lim_{π} βX X assigns to an ultrafilter its unique limit point.

Let Pre be the category of preordered sets and preorder preserving maps. The extensions of the monad βββ to Pre and Top are well studied (see, e.g., [HST14]).

We first consider the extension of βββ to Pre. If is a preorder on a set X, then we extend it to βX as follows:

χ \overline{}ξ iff S χ T ξ S T .

Here S x X s S with s x and S is defined dually. The extension \overline{}can equivalently be defined as follows:

χ \overline{}ξ iff S χ T ξ S T .

This in turn is equivalent to:

χ \overline{}ξ iff S S ξ S χ .

The last condition shows that the extension of to βX is the well-known ultrafilter extension in modal logic (see, e.g., [BRV01, Sec. 2.5]). It is straightforward to see that \overline{}is a preorder on βX. Moreover, if f X X^{}is preorder preserving, then so is βf βX βX^{}. We thus obtain an extension β̄ Pre Pre of the functor β to the category of preorders.

To extend further the monad structure, we must show that for a preorder X, the maps e_{X} X βX and m_{X} ββX βX preserve the corresponding preorders. This however holds more generally for functors satisfying the Beck-Chevalley condition (see [HST14, Thm. III.1.12.1]); that the latter condition is satisfied by β is proved in [HST14, Ex. III.1.12.3(3)].

We thus obtain a monad β̄ββ on Pre. The structure of a β̄ββ-algebra on a preorder X amounts to a preorder preserving map lim_{π} βX X, which by Manes’ theorem gives a compact Hausdorff topology π on X. As is discussed in [HST14, Ex. III.5.2.1(3)], the map lim_{π} is preorder preserving precisely if the preorder is π-closed in the product X X. Furthermore, morphisms of such algebras are simply continuous preorder preserving maps. This yields the following theorem:

4.1. Theorem. (see, e.g., [HST14, Ex. III.5.2.1(3)]) The category Pre β̄ββ of β̄ββ-algebras on the category of preorders is concretely isomorphic to PNach .

We next turn to the extension of the monad βββ to Top.

4.2. Definition. ([Sal00]) For a topology τ on a set X let τ̄ be the topology on βX given by the basis φ U U τ , where φ X βX is the Stone map. For a topological space X, τ , let β̄ X, τ βX, τ̄ .

It is easy to check that if f X, τ X^{}, τ^{}is continuous, then so is βf βX, τ̄ βX^{}, τ̄^{}. Therefore, we obtain an endofunctor β̄ Top Top. As explained in [HST14, Sec. III.5.6], the maps e_{X} X, τ βX, τ̄ and m_{X} ββX, ¯ τ̄ βX, τ̄ are continuous, and so the endofunctor β̄ extends to a monad β̄ββ on Top.

In order to obtain an analogue of Theorem 4.1 for Topβ̄ββ, we require some general definitions and facts from [HST14], together with formulations of the particular cases in the topological setting.

4.3. Definition. A preorder-enriched category is a pair C, , where C is a category and for each pair of objects X, Y in C, the hom-set C X, Y is equipped with a preorder so that g g^{}implies g f g^{}f and h g h g^{}for all f X^{}X, g, g^{}X Y, and h Y Y^{}.

4.4. Notation. For brevity, we will abuse notation and call preorder-enriched categories simply preordered categories.

We will be concerned with two examples of preordered categories. First, we view Pre as a preordered category by equipping the sets Pre X, , X^{}, ^{}with pointwise preorders; that is, for preorder preserving maps f, g X X^{}, we define f g provided f x ^{}g x for all x X. Secondly, Top can be viewed as a preordered category as follows. For f, g Top X, τ , X^{}, τ^{}, define f g provided f x _{τ^{}}g x for all x X, where _{τ^{}}is the specialization preorder of X^{}, τ^{}.

4.5. Definition. For preordered categories C, , C^{}, ^{}, a functor F C C^{}is preorder preserving provided for all objects X, Y of C, the map F_{X,Y} C X, Y C^{}F X, F Y is preorder preserving. A monad on a preordered category is preorder preserving provided its underlying endofunctor is preorder preserving.

As follows from [HST14, Sec. III.5.1, III.5.4, and III.5.6], the monad β̄ββ is preorder preserving on both Pre and Top.

4.6. Definition. A preorder preserving monad T T, e, m on a preordered category C, is of Kock-Zöberlein type ( KZ-monad) if for every object X the morphism m_{X} T T X T X is left adjoint to e_{T X} T X T T X. Since m_{X} e_{T X} is the identity morphism on T X, this amounts to requiring the inequality Id_{T T X} e_{T X} m_{X} in each C T T X, T T X . Dually, a co-KZ-monad is the one for which m_{X} is right adjoint to e_{T X}, i. e. the opposite inequality e_{T X} m_{X} Id_{T T X} holds for all objects X.

This is a particular case of [Koc95], where (among many other things) it is proved that for a KZ-monad (resp., co-KZ-monad) T on a preordered category C, if a morphism a T X X defines a T-algebra structure on X, then it is a left (resp., right) adjoint retraction for e_{X} X T X. In particular, any two T-algebra structures a_{1}, a_{2} T X X on an object X are -equivalent (meaning that a_{1} a_{2} a_{1} in C T X, X ).

It follows from [HST14, Thm. III.5.4.1] that the monad β̄ββ on Top is a co-KZ-monad (the preorder used in [HST14] is opposite to the preorder we use). In particular, if a topological space X, τ admits a β̄ββ-algebra structure, then e_{X} X β̄X has a right adjoint (with respect to the specialization preorders on X and βX). Spaces with this property are called representable in [HST14]. It follows from [HST14, Sec. III.5.7] that representable spaces have the same features as stably compact spaces, with the only exception that in general they are not T_{0}-spaces.

4.7. Definition. We call a topological space X quasi stably compact if it is compact, locally compact, stable, and sober. Let QStKSp be the category of quasi stably compact spaces and proper maps.

4.8. Remark. In presence of local compactness, the conditions of being compact, stable, and sober can be replaced by a single condition of being supersober, where we recall (see [GHKLMS03, Def. VI-6.12] or [Sal00]) that a topological space X, τ is supersober or strongly sober provided for each ultrafilter on X, the set cl_{τ} A A of limit points of is the closure of a point of X.^{1}

4.9. Remark. Clearly a quasi stably compact space is stably compact iff it is a T_{0}- space. In fact, a topological space X is quasi stably compact iff its T_{0}-reflection X_{0} is stably compact, where we recall that the T_{0}-reflection is the quotient space of X by the equivalence relation given by x y iff cl x cl y . This can be seen by observing that X has property iff X_{0} has property , where is being compact, locally compact, stable, or sober.

4.10. Remark. In view of the previous remarks, it would be natural to use the term ‘stably compact’ in the general case, avoiding the prefix ‘quasi.’ This is what Salbany does in [Sal00, p. 483]. In the T_{0}-case we could then use the term ‘stably compact T_{0}- space.’ However, since it is already established to assume T_{0} in the definition of stably compact, we opted to add the prefix ‘quasi’ in the non-T_{0}-case.

By [HST14, Thm. III.5.7.2], a topological space is quasi stably compact iff it is representable. We recall that a continuous map f X, τ X^{}, τ^{}between representable spaces is a pseudo-homomorphism if for some adjoints α of e_{X} X βX and α^{}of e_{X^{}}X^{}βX^{}, there is a -equivalence f α α^{}βf. As pointed out in [HST14, Def. III.5.4.3(2)], this condition does not depend on the particular choice of adjoints α and α^{}(i.e. replacing “some” by “any” above would give an equivalent definition). By [HST14, Prop. III.5.7.6], a continuous map between representable spaces is a pseudohomomorphism iff it is a proper map. Consequently, we obtain the following theorem:

4.11. Theorem. ([HST14, Sec. III.5.7]) The category of representable spaces and pseudohomomorphisms is concretely isomorphic to QStKSp .

As we saw, if we have a β̄ββ-algebra structure on a topological space X, τ , then the topology is a quasi stably compact topology. But more is true.

4.12. Lemma. If a topological space X, τ has a β̄ββ-algebra structure, then X, τ admits a stronger compact Hausdorff topology π such that the following conditions are satisfied:

1 X, τ is a quasi stably compact space;

2 X, π is a compact Hausdorff space;

3 τ π;

4 compact saturated sets in X, τ are closed in X, π .

1We note that in [GHKLMS03, Def. VI-6.12] it is assumed that the point is unique, and hence the space is T_{0}. We do not assume uniqueness because we do not have the T_{0} separation axiom.

Proof. The structure of a β̄ββ-algebra on a topological space X, τ gives in particular the structure of a βββ-algebra on the set X, so applying Manes’ theorem yields a compact Hausdorff topology π on X, determined by the map lim_{π} βX X. That this is not only a βββ-algebra structure but also a β̄ββ-algebra structure means that lim_{π} βX, τ̄ X, τ is continuous. Since β̄ββ is a co-KZ-monad, lim_{π} is a right adjoint to e_{X,τ} X, τ βX, τ̄ . As noted above, this implies that X, τ is a representable space, hence a quasi stably compact space. Therefore, Conditions (1) and (2) are satisfied.

To verify Condition (3), observe that it is equivalent to showing that lim_{π} χ is a τ-limit point of χ for any ultrafilter χ βX. This amounts to showing that any τneighborhood of lim_{π} χ belongs to χ; in other words, that e_{X} lim_{π} χ τ χ. But this means that e_{X} lim_{π} χ τ̄ χ, where τ̄ is the specialization preorder of τ̄, which holds for any χ since lim_{π} is adjoint to e_{X}.

It remains to verify Condition (4). Observe that m_{X} lim_{Π} β̄β̄ X, τ β̄ X, τ equips βX, τ̄ with a β̄ββ-algebra structure, and lim_{π} is a homomorphism of β̄ββ-algebras. In particular, lim_{π} is a homomorphism of βββ-algebras, and hence lim_{π} βX, Π X, π is a continuous map. On the other hand, since lim_{π} m_{X} lim_{π} β lim_{π} , m_{X} is adjoint to e_{βX,τ̄}, and lim_{π} is adjoint to e_{X,τ}, we conclude that lim_{π} is a pseudo-homomorphism, hence lim_{π} βX, τ̄ X, τ is a proper map. Now, if K is a compact saturated set in X, τ , then lim_{π}^{1} K is compact saturated in βX, τ̄ since lim_{π} is proper. By [BMM08, Thm. 2.12], τ̄ is the intersection of Π and the Alexandroff topology of -upsets, where we recall from Section 3 that χ ξ iff χ τ ξ. Therefore, lim_{π}^{1} K is a Π-closed -upset. As lim_{π} is an onto continuous map between compact Hausdorff spaces βX, Π and X, π , it follows that lim_{π} lim_{π}^{1} K K is π-closed.

Lemma 4.12 motivates the following definition.

4.13. Definition. Let X, τ, π be a bitopological space. We call X, τ, π a bitopological quasi stably compact space if

1 X, τ is a quasi stably compact space;

2 X, π is a compact Hausdorff space;

3 τ π;

4 compact saturated sets in X, τ are closed in X, π .

Let BQStKSp be the category of bitopological quasi stably compact spaces and bicontinuous maps.

4.14. Remark. If X, τ, π and X^{}, τ^{}, π^{}are bitopological quasi stably compact spaces and f X, τ, π X^{}, τ^{}, π^{}is bicontinuous, then f X, τ X^{}, τ^{}is proper. To see this, let K be compact saturated in X^{}, τ^{}. Then K is closed in X^{}, π^{}. Since f X, π X^{}, π^{}is continuous, f^{1} K is closed in X, π . Therefore, f^{1} K is compact in X, π , hence compact in X, τ . Thus, f is proper.

4.15. Theorem. The category Topβ̄ββ of algebras over the monad β̄ββ on Top is concretely isomorphic to BQStKSp .

Proof. By Lemma 4.12, if a topological space X, τ has a β̄ββ-algebra structure, then X, τ admits a stronger compact Hausdorff topology π such that X, τ, π is a bitopological quasi stably compact space.

Conversely, suppose X, τ, π is a bitopological quasi stably compact space. Since X, π is compact Hausdorff, by Manes’ theorem, lim_{π} βX, Π X, π is continuous. From τ π it follows that e_{X} lim_{π} χ τ̄ χ. To see that lim_{π} is a right adjoint to e_{X,τ}, we must show that lim_{π} is a preorder preserving map from βX, \overline{_{τ}} to X, _{τ} , where _{τ} is the specialization preorder of τ. This in turn follows from continuity of lim_{π} βX, \overline{τ} X, τ . To see the latter, let χ βX. Set x lim_{π} χ . Suppose U is a τopen neighborhood of x. Since X, τ is locally compact, there are an open neighborhood V of x and a compact set K with V K U. Since the saturation of K is compact and is contained in U, without loss of generality we may assume that K is compact saturated. We set U^{}φ V . Then x lim_{π} χ V implies V χ, so χ φ V U^{}, and hence U^{}is a τ̄-neighborhood of χ. To see that lim_{π} U^{}U, let x^{}lim_{π} U^{}. Then there is χ^{}U^{}φ V with lim_{π} χ^{}x^{}. Therefore, V χ^{}. But lim_{π} χ^{}x^{} yields that x^{}belongs to the π-closure of V. Since K is compact saturated in X, τ , by Condition (4), it is π-closed. Thus, x^{}K U, and hence lim_{π} U^{}U. Consequently, lim_{π} βX, τ̄ X, τ is continuous. This gives that lim_{π} is a right adjoint to e_{X}, and so lim_{π} equips X, τ with a β̄ββ-algebra structure.

Finally, a β̄ββ-algebra homomorphism from X, τ, lim_{π} to X^{}, τ^{}, lim_{π^{}}is a continuous map from X, τ to X^{}, τ^{}compatible with lim_{π} and lim_{π^{}}. Therefore, it is also a continuous map from X, π to X^{}, π^{}. Such homomorphisms are precisely the bicontinuous maps between bitopological quasi stably compact spaces. Thus, we indeed obtain the required isomorphism of categories.

4.16. Corollary. The categories BQStKSp and PNach are concretely isomorphic.

Proof. By [HST14, Cor. III.5.4.2], Pre β̄ββ and Topβ̄ββ are concretely isomorphic. By Theorem 4.1, Pre β̄ββ is concretely isomorphic to PNach . By Theorem 4.15, Topβ̄ββ is concretely isomorphic to BQStKSp . The result follows.

5. More on the isomorphism of BQStKSp and PNach

In the previous section we deduced the isomorphism of BQStKSp and PNach from a general result in [HST14, Cor. III.5.4.2]. But this isomorphism can also be seen as a direct generalization of the isomorphism of StKSp and Nach [GHKLMS03, Prop. VI-6.23]. In this section we give the details of how such a generalization works.

5.1. Lemma. Let X, τ, π be a bitopological quasi stably compact space and let be the specialization preorder of τ. Then X, π, is a preordered Nachbin space.

Proof. Suppose x y. Then there is W τ with x W and y W. As X, τ is locally compact, there are an open set U and a compact set K in X, τ such that x U K W. Since the saturation of K is compact and is contained in W, without loss of generality we may assume that K is compact saturated, hence closed in X, π . Let V X K. Then x U, y V, U is an open upset, V is an open downset in X, π, , and U V . Therefore, U V is an open neighborhood of x, y missing . Thus, is closed, and hence X, π, is a preordered Nachbin space.

5.2. Lemma. Suppose X, π, is a preordered Nachbin space.

1 The specialization preorder of π^{u} coincides with .

2 Compact saturated subsets of X, π^{u} are exactly the closed upsets of X, π, .

3 Compact saturated subsets of X, π^{d} are exactly the closed downsets of X, π, .

Proof. (1) Since is closed in X X, it is clear that x y iff for each open upset U, we have x U implies y U. Thus, the specialization preorder of π^{u} coincides with .

(2) If K is a closed upset in X, π, , then K is compact in X, π , hence compact in X, π^{u} . By (1), it is also saturated in X, π^{u} . Therefore, K is compact saturated in X, π^{u} . Conversely, suppose that K is compact saturated in X, π^{u} . By (1), K is an upset in X, π, . Let x K. Then y x for each y K. Since is closed in X X, for each such y, there are an open upset U_{y} and an open downset V_{y} such that y U_{y}, x V_{y}, and U_{y} V_{y} . The U_{y} provide an open cover of K in X, π^{u} . Therefore, there is a finite subcover U_{y_{1}}, …, U_{y_{n}}. Set U U_{y_{1}} U_{y_{n}} and V V_{y_{1}} V_{y_{n}}. Then K U, U V , and x V. Thus, x V X K. This yields that X K is open, hence K is closed in X, π .

(3) If X, π, is a preordered Nachbin space, then so is X, π, ^{op} . Now apply (2).

Let X, be a preorder. For S X, we call m S a maximal point of S if m s and s S imply s m; minimal points are defined dually. Let max S be the set of maximal points and min S the set of minimal points of S. We call S a chain provided s t or t s for all s, t S. The next lemma generalizes [Esa85, Thm. III.2.1].

5.3. Lemma. Let X, π, be a preordered Nachbin space. If F is a closed subset of X, π , then for each x F there exist m max F and l min F such that l x m.

Proof. Let x F. Since is closed in X X, by [Nac65, Prop. 1], x y X x y is closed, and hence x F is closed. Let C be a maximal chain in x F starting at x. Then c F c C is a family of closed sets with the finite intersection property. Since X, π is compact, there is m c F c C . Clearly x m, and since C is a maximal chain, it follows that m max F . That there is l min F with l x is proved similarly.

5.4. Lemma. If X, π, is a preordered Nachbin space, then X, π^{u} and X, π^{d} are quasi stably compact spaces.

Proof. We prove that X, π^{u} is a quasi stably compact space. That X, π^{d} is quasi stably compact is proved by switching to ^{op}. Clearly X, π^{u} is compact. By Lemma 5.2(2), compact saturated sets in X, π^{u} are closed upsets in X, π, . From this it is immediate that X, π^{u} is stable. It is well known that in a preordered Nachbin space, if U π^{u}, then U V π^{u} cl_{π} V U , where for A X, we recall that A x X a x for some a A . Since cl_{π} V is a closed upset (see [Nac65, Prop. 4] or [BMM02, Prop. 2.3]), it is compact saturated in X, π^{u} by Lemma 5.2(2). This yields that X, π^{u} is locally compact. It remains to show that X, π^{u} is sober.

Let F be an irreducible closed set in X, π^{u} . Since F is closed in X, π^{u} , we see that F is a closed downset in X, π, . By Lemma 5.3, for each x F, there is m max F with x m. We show that F m for some (and hence all) m max F . If not, then for each m max F there is n max F with n m. As is closed in X X, there are an open upset U_{n} and an open downset V_{m} such that n U_{n}, m V_{m}, and U_{n} V_{m} . This yields U_{n} cl_{π} V_{m} . The V_{m} cover F, so by compactness, there are finitely many V_{m} covering F. Therefore, finitely many cl_{π} V_{m} cover F. Since the cl_{π} V_{m} are closed in X, π^{u} and F is irreducible, there is one m max F with F cl_{π} V_{m} . The obtained contradiction proves that F m for each m max F . Thus, F cl_{π^{u}} m , and hence X, π^{u} is sober.

5.5. Theorem. The categories BQStKSp and PNach are concretely isomorphic.

Proof. Define a functor F BQStKSp PNach as follows. If X, τ, π is a bitopological quasi stably compact space, then set F X, τ, π X, π, , where is the specialization preorder of τ; and if f X, τ, π X^{}, τ^{}, π^{}is a bicontinuous map between bitopological quasi stably compact spaces, then set F f f. By Lemma 5.1, F X, τ, π is a preordered Nachbin space. It is obvious that F f is continuous and preorder preserving. Thus, F is well-defined.

Define a functor G PNach BQStKSp as follows. If X, π, is a preordered Nachbin space, then set G X, π, X, π^{u}, π ; and if f X, π, X^{}, π^{}, ^{}is a continuous preorder preserving map, then set G f f. By Lemmas 5.2 and 5.4, G X, π, is a bitopological quasi stably compact space. It is also obvious that G f is bicontinuous. Thus, G is well-defined.

We show that if X, τ, π is a bitopological quasi stably compact space, then τ π^{u}. Since τ π and is the specialization preorder of τ, it is clear that τ π^{u}. Conversely, suppose U π^{u}. Let x U. Then for each y U, we have x y. Therefore, there is U_{y} τ with x U_{y} and y U_{y}. Since X, τ is locally compact, there are open sets V_{y} and compact saturated sets K_{y} in X, τ such that x V_{y} K_{y} U_{y}. Clearly K_{y} y U U^{c} . As each K_{y} is closed in X, π and X, π is compact, there are V_{y_{1}}, …, V_{y_{n}} and K_{y_{1}}, …, K_{y_{n}} with x V_{y_{1}} V_{y_{n}} K_{y_{1}} K_{y_{n}} U. Thus, we found an open neighborhood V V_{y_{1}} V_{y_{n}} of x in X, τ contained in U, so U τ, which completes the proof.

It follows that for a bitopological quasi stably compact space X, τ, π , we have GF X, τ, π X, τ, π . It also follows from Lemma 5.2(1) that if X, π, is a preordered Nachbin space, then the specialization preorder of π^{u} coincides with , so F G X, π, X, π, . Consequently, the functors F and G establish a concrete isomorphism of BQStKSp and PNach .

5.6. Remark. If X, π, is a Nachbin space, then the corresponding bitopological quasi stably compact space X, π^{u}, π yields a stably compact space X, π^{u} . Moreover, π is uniquely determined from π^{u} as the patch topology. Conversely, if X, τ is stably compact and π is the patch topology, then X, τ, π is a bitopological quasi stably compact space, and the corresponding preordered Nachbin space X, π, is a Nachbin space since X, τ is a T_{0}-space, so is a partial order. Thus, the isomorphism of Theorem 5.5 restricts to the well-known isomorphism between StKSp and Nach [GHKLMS03, Prop. VI-6.23].

6. Quasi spectral spaces

Spectral spaces were generalized to quasi spectral spaces in [BMM08]. A topological space is quasi spectral if it is a coherent supersober space. Equivalently, X is quasi spectral provided X is compact, coherent, and sober. Consequently, quasi spectral spaces generalize spectral spaces by dropping the T_{0}-separation axiom. Not surprisingly, a topological space X is quasi spectral iff its T_{0}-reflection is spectral [BMM08, Thm. 4.6].

Let QSpec be the category of quasi spectral spaces and spectral maps. Since each coherent space is locally compact, each quasi spectral space is quasi stably compact. Moreover, the same argument as in the case of spectral spaces gives that a continuous map between quasi spectral spaces is proper iff it is spectral. Therefore, QSpec is a full subcategory of QStKSp .

An important example of a quasi spectral space is given by βX, τ̄ from Definition 4.2. Indeed, as follows from [Sal00, Sec. 2] (see also [BMM08, Prop. 4.2]), for an arbitrary topological space X, τ , the space βX, τ̄ is quasi spectral and e_{X} X βX is an embedding. By [Sal00, Prop. 3], if X, τ is quasi stably compact, then there is a (not necessarily unique) retraction r_{X} βX X such that r_{X} e_{X} Id_{X}. Consequently, quasi stably compact spaces are precisely the retracts of quasi spectral spaces. This provides a generalization of the well-known characterization of stably compact spaces as retracts of spectral spaces; see [Joh82, Thm. VII.4.6] or [Sim82, Lem. 3.13].

We next introduce bitopological analogues of quasi spectral spaces.

6.1. Definition. Let X, τ, π be a bitopological space. We call X, τ, π a bitopological quasi spectral space if

1 X, τ is a quasi spectral space;

2 X, π is a Stone space;

3 τ π; 4 compact opens in X, τ are clopen in X, π .

Let BQSpec be the category of bitopological quasi spectral spaces and bicontinuous maps.

6.2. Remark. If X, τ, π and X^{}, τ^{}, π^{}are bitopological quasi spectral spaces and f X, τ, π X^{}, τ^{}, π^{}is bicontinuous, then f X, τ X^{}, τ^{}is spectral. To see this, let U be compact open in X^{}, τ^{}. Then U is clopen in X^{}, π^{}. Since f X, π X^{}, π^{}is continuous, f^{1} U is clopen in X, π . Therefore, f^{1} U is compact in X, π , hence compact in X, τ . Thus, f is spectral.

6.3. Remark. The category BQSpec is a full subcategory of BQStKSp . To see this it is sufficient to observe that each bitopological quasi spectral space is a bitopological quasi stably compact space. Let X, τ, π be a bitopological quasi spectral space. Since each quasi spectral space is quasi stably compact, X, τ is quasi stably compact. It is also clear that X, π is compact Hausdorff and τ π. It is well known that in a spectral space compact saturated sets are intersections of compact opens. The same is true in quasi spectral spaces. Therefore, since compact opens in X, τ are clopen in X, π , compact saturated sets in X, τ are closed in X, π . Thus, X, τ, π is a bitopological quasi stably compact space.

We next show that the isomorphism between BQStKSp and PNach restricts to an isomorphism of BQSpec and PPries .

6.4. Lemma. Let X, τ, π be a bitopological quasi spectral space and let be the specialization preorder of τ. Then X, π, is a preordered Priestley space.

Proof. If X, τ, π is a bitopological quasi spectral space, then X, π is a Stone space. Moreover, if x y, then there is a compact open U in X, τ with x U and y U. Therefore, U is a clopen upset in X, π, containing x but not y. Thus, X, π, is a preordered Priestley space.

6.5. Lemma. If X, π, is a preordered Priestley space, then compact opens of X, π^{u} are exactly the clopen upsets and compact opens of X, π^{d} are exactly the clopen downsets of X, π, .

Proof. We prove that compact opens in X, π^{u} are clopen upsets in X, π, . That compact opens in X, π^{d} are clopen downsets is proved by switching to ^{op}. Let U be a clopen upset in X, π, . Then U is open in X, π^{u} . Moreover, U is compact in X, π , which makes it compact in X, π^{u} . Thus, U is compact open in X, π^{u} . Conversely, let U be compact open in X, π^{u} . Then it is open in X, π and an upset. Take y U. For each x U, since U is an upset, x y. By the Priestley separation axiom, there is a clopen upset V_{x} with x V_{x} and y V_{x}. The V_{x} cover U and are open in X, π^{u} . Therefore, compactness of U implies that there are x_{1}, …, x_{n} U with U V_{x_{1}} V_{x_{n}} V_{y}. Thus, we have a clopen upset V_{y} containing U and missing y. This yields U^{c} V_{y} y U . Since U^{c} is closed in X, π and X, π is compact, there are y_{1}, …, y_{m} U with U^{c} V_{y_{1}} V_{y_{m}} . If W is the intersection of the V_{y_{i}}, then W is a clopen upset, W contains U, and W misses U^{c}. Thus, W U, and hence U is a clopen upset.

6.6. Lemma. If X, π, is a preordered Priestley space, then X, π^{u} and X, π^{d} are quasi spectral spaces.

Proof. Clearly X, π^{u} is compact. By Lemma 5.4, X, π^{u} is sober. By Lemma 6.5, compact opens of X, π^{u} are clopen upsets of X, π, . Since in a preordered Priestley space, each open upset is the union of clopen upsets contained in it and finite intersections of clopen upsets are clopen upsets, X, π^{u} is coherent. Thus, X, π^{u} is a quasi spectral space. That X, π^{d} is also quasi spectral is proved by switching to ^{op}.

6.7. Theorem. The categories BQSpec and PPries are concretely isomorphic.

Proof. Let F BQStKSp PNach and G PNach BQStKSp be the functors from Theorem 5.5. By Lemmas 6.4–6.6, F restricts to a functor BQSpec PPries and G restricts to a functor PPries BQSpec . Thus, by Theorem 5.5, the restrictions yield a concrete isomorphism of BQSpec and PPries .

Thus, we obtain the following commutative diagram, which generalizes the commutative diagram of Section 2.

BQStKSp  PNach

BQSpec  PPries

6.8. Remark. Theorem 6.7 generalizes the well-known isomorphism of Spec and Pries [Cor75].

6.9. Remark. An alternate approach to Theorem 6.7 was pointed out to us by the referee. Let X, π, be a preordered Nachbin space. The cone f_{i} X X_{i} of PNach - morphisms from X, π, to finite PNach -objects X_{i}, _{i} is initial provided both π and are determined by the cone; that is, π is generated by f_{i}^{1} U , where U X_{i}, and x y iff f_{i} x _{i} f_{i} y for all f_{i} in the cone.^{2} It is immediate from the definition of a preordered Priestley space that the cone f_{i} X X_{i} is initial iff X, π, is a preordered Priestley space.

Similarly, for X, τ, π BQStKSp , the cone f_{i} X X_{i} of BQStKSp -morphisms from X, τ, π to finite BQStKSp -objects is initial (that is, both τ and π are determined by the cone) iff X, τ, π BQSpec .

To obtain Theorem 6.7 it now suffices to observe that the concrete isomorphism between BQStKSp and PNach of Corollary 4.16 preserves and reflects initial cones, and restricts to an isomorphism between the subcategories of finite objects of BQStKSp and PNach , respectively. Therefore, it carries bitopological quasi spectral spaces to preordered Priestley spaces and vice versa. Thus, it restricts to an isomorphism between BQSpec and PPries .

2Since X_{i} is finite, π_{i} is discrete.

This approach also permits to deduce Lemma 6.6 from a result of [HN14]. Similarly to the above, an object X, τ of QStKSp is in QSpec iff the cone f_{i} X X_{i} of QStKSp - morphisms from X, τ to finite QStKSp -objects (that is, finite topological spaces) is initial. Now apply [HN14, Prop. 2.1] to obtain that if X, π, PPries , then K X, π, QSpec .

6.10. Remark. It follows from Remark 6.9 that the initial cones f_{i} X X_{i} in PNach and BQStKSp are in fact limiting cones. In other words, each object in PPries and BQSpec is the inverse limit of the inverse system of all its finite images. This is not true in QStKSp or Top.

For a simple example in QStKSp , let X be an infinite set with the trivial topology τ. Then X, τ QSpec . Therefore, as we will see in Theorem 6.16, X, τ is homeomorphic to the inverse limit of an inverse system of finite spaces. However, the inverse limit of all finite images of X, τ is homeomorphic to βX with the trivial topology, hence is not homeomorphic to X, τ .

For an example in Top, let X be the ordinal ω 1 and let τ be the topology of open downsets. It is easy to see that X, τ is a spectral space. In particular, X, τ is a T_{0}-space. By [Hoc69, Joy71], X, τ is homeomorphic to the inverse limit of an inverse system of finite T_{0}-spaces. In fact, X, τ is homeomorphic to the inverse limit of all finite T_{0}-spaces that are spectral images of X, τ . However, X, τ is not homeomorphic to the inverse limit of all finite T_{0}-images of X, τ . To see this, observe that finite T_{0}-quotients of X, τ are obtained by breaking X into finitely many intervals, exactly one of which is infinite. The quotient map is spectral iff the infinite interval contains ω. For example, if we break X into 0, ω and ω , then the corresponding quotient map is not spectral. Because of this, choosing the point corresponding to the infinite interval in each quotient space produces a new point in the inverse limit. In fact, the inverse limit is homeomorphic to X^{}, τ^{}, where X^{}ω 2 and τ^{}is the topology of open downsets in ω 2. Clearly the same example works also for the category of T_{0}-spaces.

The next corollary is an immediate consequence of [Dia07, Sec. 2] and Theorems 5.5 and 6.7.

6.11. Corollary. Profinite objects in BQStKSp are exactly the bitopological quasi spectral spaces.

Moreover, since the functor K PNach Top factors through BQStKSp , from Theorems 3.2 and 6.7 we conclude:

6.12. Corollary. A topological space X, τ is profinite iff it admits a stronger Stone topology π such that X, τ, π is a bitopological quasi spectral space.

Next comes our key result, that each quasi stably compact space X, τ admits a stronger compact Hausdorff topology π such that X, τ, π BQStKSp , and that if X, τ is quasi spectral, then X, τ, π BQSpec .

6.13. Lemma. Let Y, σ be a compact Hausdorff space, X be a set, and p X Y, s Y X be maps with the composite ps equal to the identity map on Y. Then there is a compact Hausdorff topology π on X such that both p and s are continuous. Moreover, for each y Y, the subspace topology on p^{1} y is the same as the one-point compactification of the discrete space p^{1} y s y . Furthermore, if Y is a Stone space, then X, π is a Stone space.

Proof. Let be the collection of subsets of X of the form p^{1} U F, where U σ, F is a finite subset of X s Y , and denotes symmetric difference of sets. Since

p^{1} U F p^{1} U^{}F^{}p^{1} U U^{}p^{1} U F^{}p^{1} U^{}F F F^{}

and p^{1} U F^{}p^{1} U^{}F F F^{}F F^{}is a finite subset of X s Y , we see that is closed under finite intersections. In addition, X p^{1} Y . Therefore, generates a topology π on X.

To see that X, π is Hausdorff, let x, x^{}X with x x^{}. If p x p x^{}, then as Y, σ is Hausdorff, there are disjoint U, U^{}σ separating p x , p x^{}, so p^{1} U , p^{1} U^{}are disjoint and separate x, x^{}. On the other hand, if p x p x^{}, then one of x, x^{}does not belong to the s-image of Y, and without loss of generality we may assume that x s Y . Thus, x , so x and X x p^{1} Y x are disjoint open sets of X, π and separate x, x^{}.

To see that X, π is compact, let p^{1} U_{i} F_{i} i I be a cover of X with elements of . We show that the U_{i} cover Y. Let y Y. Then there is i I with s y p^{1} U_{i} F_{i}. Since s y F_{i}, we see that s y p^{1} U_{i} , hence y ps y U_{i}. As Y, σ is compact, there is a finite subcover U_{i_{1}}, …, U_{i_{n}}. Therefore, p^{1} U_{i_{1}} , …, p^{1} U_{i_{n}} is a cover of X. Thus, p^{1} U_{i_{1}} F_{i_{1}} p^{1} U_{i_{n}} F_{i_{n}} misses at most finitely many points of X since it contains the complement of F_{i_{1}} F_{i_{n}}. Adding finitely many p^{1} U_{i} F_{i} will then produce a finite subcover, yielding compactness of X, π .

That p is continuous follows from the definition of π, and continuity of s follows since

s^{1} p^{1} U F s^{1}p^{1} U s^{1} F U U.

Next, let y Y. The subspace topology on p^{1} y is generated by the sets

p^{1} y p^{1} U F p^{1} y U p^{1} y F .

If y U, then such sets are cofinite in p^{1} y and contain s y ; and if y U, then such sets are finite and do not contain s y . This is precisely the topology obtained by compactifying the discrete space p^{1} y s y with the point s y .

Finally, if Y, σ is a Stone space, let _{0} be the subset of of those p^{1} U F, where U is clopen in Y, σ . Then _{0} is also closed under complements as

X p^{1} U F X p^{1} U F X p^{1} U F
X p^{1} U F p^{1} Y U F.

Thus, _{0} is a Boolean algebra. As Y, σ is a Stone space, it is clear that _{0} and generate the same topology. Hence, X, π is a Stone space.

6.14. Theorem. The forgetful functors BQStKSp QStKSp and BQSpec QSpec are surjective on objects.

Proof. Let X, τ be quasi stably compact and let X_{0}, τ_{0} be its T_{0}-reflection with the reflection map p X X_{0}. Then X_{0}, τ_{0} is stably compact. If π_{0} is the patch topology of τ_{0}, then X_{0}, π_{0} is compact Hausdorff. Choose any section s X_{0} X of p X X_{0}. By Lemma 6.13, there is a compact Hausdorff topology π on X such that p X, π X_{0}, π_{0} is continuous. We claim that X, τ, π is a bitopological quasi stably compact space. Clearly X, τ is quasi stably compact and X, π is compact Hausdorff. Let U τ. Then p U τ_{0} π_{0}, so U p^{1}p U π. Therefore, τ π. Let K be compact saturated in X, τ . Then p K is compact saturated in X_{0}, τ_{0} , so p K is closed in X_{0}, π_{0} . Thus, K p^{1}p K is closed in X, π , yielding that X, τ, π BQStKSp . Consequently, the forgetful functor BQStKSp QStKSp is surjective on objects.

If in addition X, τ QSpec , then X_{0}, τ_{0} is a spectral space, hence X_{0}, π_{0} is a Stone space. By Lemma 6.13, π is a Stone topology on X. Therefore, X, τ is a quasi spectral space, X, π is a Stone space, and τ π. Let U be compact open in X, τ . Then p K is compact open in X_{0}, τ_{0} , so p K is clopen in X_{0}, π_{0} . Thus, K p^{1}p K is clopen in X, π , yielding that X, τ, π BQSpec . Consequently, the forgetful functor BQSpec QSpec is also surjective on objects.

6.15. Remark. While the forgetful functors BQStKSp QStKSp and BQSpec QSpec are surjective on objects, neither is an equivalence as they map non-isomorphic objects to isomorphic objects. To see this, take any two Stone topologies π_{1} and π_{2} on the same set X so that X, π_{1} and X, π_{2} are not homeomorphic, and let τ be the trivial topology on X. Then X, τ, π_{1} and X, τ, π_{2} are non-isomorphic objects in BQSpec , but their images under the forgetful functor coincide with X, τ .

In fact, BQStKSp and QStKSp are not equivalent, and neither are BQSpec and QSpec . This can be seen as follows. By Theorem 4.15, BQStKSp is isomorphic to Topβ̄ββ, hence is complete. On the other hand, it is easy to see that QStKSp does not have equalizers, and the same is true about BQSpec and QSpec .^{3} To see this, let X ℕ _{1}, _{2} be the “double limit” space, where all points in ℕ are isolated, while the neighborhoods of _{i} are cofinite subsets of X containing both _{1} and _{2}. Clearly the T_{0}-reflection of X is homeomorphic to the one-point compactification of ℕ, hence is a Stone space. In particular, it is a spectral space, and so X is a quasi spectral space. Let f X X be the identity on ℕ and switch _{1} and _{2}. It is easy to see that f is a homeomorphism, hence a spectral map. If f and Id_{X} have an equalizer g X^{}X, where X^{}QStKSp and g is a QStKSp -morphism, then g X^{}is a compact subset of X missing _{1} and _{2}. Therefore, g X^{}is a finite subset of ℕ. But any finite subset of X _{1}, _{2} also belongs to QStKSp and equalizes f and Id_{X}. Therefore, not every QStKSp -morphism equalizing f and Id_{X} factors through g, so there is no equalizer for Id_{X} and f in QStKSp .

3We would like to thank the referee for simplifying our original argument considerably.

Putting Corollary 6.12 and Theorem 6.14 together yields the main result of the paper.

6.16. Theorem. For a topological space X, τ , the following are equivalent:

1 X, τ is a profinite topological space.

2 X, τ is a quasi spectral space.

3 X, τ admits a stronger Stone topology π such that X, τ, π is a bitopological quasi spectral space.

The well-known theorem [Hoc69, Joy71] characterizing profinite T_{0}-spaces is now an immediate consequence of Theorem 6.16.

6.17. Corollary. ([Hoc69, Joy71]) A T_{0}-space X, τ is profinite iff X, τ is a spectral space.

Since X, τ is quasi spectral iff its T_{0}-reflection is spectral, we conclude:

6.18. Corollary. A topological space is profinite iff its T_{0}-reflection is a profinite T_{0}- space.

References

G. Bezhanishvili. Lattice subordinations and Priestley duality. Algebra Universalis, 70(4):359–377, 2013.

G. Bezhanishvili, R. Mines, and P. J. Morandi. The Priestley separation axiom for scattered spaces. Order, 19(1):1–10, 2002.

G. Bezhanishvili, R. Mines, and P. J. Morandi. Topo-canonical completions of closure algebras and Heyting algebras. Algebra Universalis, 58(1):1–34, 2008.

P. Blackburn, M. de Rijke, and Y. Venema. Modal Logic. Cambridge University Press, 2001.

W. H. Cornish. On H. Priestley’s dual of the category of bounded distributive lattices. Mat. Vesnik, 12(27)(4):329–332, 1975.

M. Dias. Priestley spaces: the threefold way. Pré-Publicações do Departamento de Matemática 07-25, University of Coimbra 2007, 9 pages, electronic version available

    at http://www.mat.uc.pt/preprints/ps/p0725.pdf.

L. L. Esakia. Heyting Algebras I. Duality Theory (Russian). “Metsniereba”, Tbilisi, 1985.

G. Gierz, K. H. Hofmann, K. Keimel, J. D. Lawson, M. Mislove, and D. S. Scott. Continuous lattices and domains, volume 93 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, 2003.

M. Hochster. Prime ideal structure in commutative rings. Trans. Amer. Math. Soc., 142:43–60, 1969.

D. Hofmann. On a generalization of the Stone-Weierstrass theorem. Appl. Categ. Structures, 10(6):569–592, 2002.

D. Hofmann and P. Nora. Dualities in modal logic from the point of view of triples. Algebra Universalis, 73(3):297–320, 2015.

D. Hofmann, G. J. Seal, and W. Tholen. Lax algebras. In Monoidal topology. A categorical approach to order, metric and topology, volume 153 of Encyclopedia of Mathematics and its Applications, pages 145–283. Cambridge University Press, Cambridge, 2014.

G. Janelidze and M. Sobral. Profinite relational structures. Cah. Topol. Géom. Différ. Catég., 49(4):280–288, 2008.

P. T. Johnstone. Stone spaces. Cambridge University Press, Cambridge, 1982.

A. Joyal. Spectral spaces and distributive lattices. Notices Amer. Math. Soc., 18:393–394, 1971.

A. Kock. Monads for which structures are adjoint to units. J. Pure Appl. Algebra, 104(1):41–59, 1995.

E. Manes. A triple theoretic construction of compact algebras. In Sem. on Triples and Categorical Homology Theory (ETH, Zürich, 1966/67), pages 91–118. Springer, Berlin, 1969.

L. Nachbin. Topology and order. D. Van Nostrand Co., Inc., Princeton, N.J.-Toronto, Ont.-London, 1965.

H. A. Priestley. Representation of distributive lattices by means of ordered Stone spaces. Bull. London Math. Soc., 2:186–190, 1970.

S. Salbany. Ultrafilter spaces and compactifications. Portugal. Math., 57(4):481–492, 2000.

H. Simmons. A couple of triples. Topology Appl., 13(2):201–223, 1982.

T. P. Speed. Profinite posets. Bull. Austral. Math. Soc., 6:177–183, 1972.

W. Tholen. Ordered topological structures. Topology Appl., 156(12):2148–2157, 2009.

Department of Mathematical Sciences, New Mexico State University, Las Cruces NM 88003

A. Razmadze Mathematical Institute, Tbilisi State University, University St. 2, Tbilisi 0186, Georgia

A. Razmadze Mathematical Institute, Tbilisi State University, University St. 2, Tbilisi 0186, Georgia

Department of Mathematical Sciences, New Mexico State University, Las Cruces NM 88003

Email: guram@nmsu.edu, gabelaia@gmail.com, jib@rmi.ge, pmorandi@nmsu.edu
This article may be accessed at http://www.tac.mta.ca/tac/

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available from the journal’s server at http://www.tac.mta.ca/tac/. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e is required. Articles in PDF format may be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com Valeria de Paiva: Nuance Communications Inc: valeria.depaiva@gmail.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Radboud University Nijmegen: i.moerdijk@math.ru.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Ljubljana: Alex.Simpson@fmf.uni-lj.si James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Université du Québec à Montréal : tierney.myles4@gmail.com R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca